# New Dataset and Methods for Fine-Grained Compositional Referring Expression Comprehension via Specialist-MLLM Collaboration

### v2

@@ Abstract
Referring Expression Comprehension (REC) is a foundational cross-modal task that evaluates the interplay of language understanding, image comprehension, and language-to-image grounding. It serves as an essential testing ground for Multimodal Large Language Models (MLLMs). To advance this field, we introduce a new REC dataset characterized by two key features. First, it is designed with controllable difficulty levels, requiring multi-level fine-grained reasoning across object categories, attributes, and multi-hop relationships. Second, it incorporates negative text and images generated through fine-grained editing and augmentation, explicitly testing a model’s ability to reject scenarios where the target object is absent—an often-overlooked yet critical challenge in existing datasets. To address fine-grained compositional REC, we propose novel methods based on a Specialist-MLLM collaboration framework, leveraging the complementary strengths of these models: Specialist Models excel at simpler tasks with high efficiency, while MLLMs are better suited for complex reasoning. Building on this synergy, we introduce two collaborative strategies. The first, Slow-Fast Adaptation (SFA), employs a routing mechanism to adaptively delegate simple tasks to Specialist Models and complex tasks to MLLMs. Additionally, common error patterns in both models are mitigated through a simple yet effective target-refocus strategy. The second, Candidate Region Selection (CRS), generates multiple bounding box candidates for the target object based on Specialist Model and uses the advanced reasoning capabilities of MLLMs to identify the correct target. Extensive experiments on our proposed dataset and other challenging compositional REC benchmarks validate the effectiveness of our approaches. The SFA strategy achieves a superior trade-off between localization accuracy and efficiency, and the CRS strategy significantly boosts the performance of both Specialist Models and MLLMs. We aim for this work to offer valuable insights into solving complex real-world tasks by strategically combining existing tools for maximum effectiveness, rather than reinventing them.
@@ I Introduction
To address these limitations, we introduce FineCops-Ref, a benchmark specifically designed to enhance fine-grained reasoning in REC. Our dataset incorporates controlled difficulty levels, compelling MLLMs to reason across object categories, attributes, and multi-hop relationships. The difficulty levels are categorized based on the number of attributes and relationships necessary for accurately localizing the target object. Specifically, if an image contains only one possible target, the difficulty level is set to 1, regardless of the complexity of the expression. When there are multiple objects of the same category as the target object in the image, and they can be distinguished by a single attribute or relationship, the difficulty level is assigned as 2. Conversely, if the model is required to interpret two or more relationships and attribute dependencies, the difficulty level is set to 3. Beyond compositional complexity, FineCops-Ref introduces negative samples carefully constructed through meticulous editing. These samples evaluate model resilience against misalignments and hallucinations, providing a more rigorous assessment of their true visual grounding capabilities. By addressing both compositional reasoning and negative sample inclusion, FineCops-Ref establishes a more realistic and challenging benchmark for advancing REC research in the MLLM era.
@@
While MLLMs demonstrate strong generalization and reasoning abilities, they often struggle with fine-grained localization and compositional grounding, particularly when precise object differentiation is required. In contrast, Specialist Models, which are typically lightweight and trained for specific vision tasks, excel in efficient and accurate low-level perception but lack the broad contextual reasoning capabilities of MLLMs. Given their complementary strengths, an effective solution is to leverage both models strategically—assigning simpler, perception-driven tasks to Specialist Models while reserving MLLMs for complex reasoning. This collaborative approach not only enhances accuracy and robustness but may also improve computational efficiency, as MLLMs are invoked only when necessary.
@@
Motivated by this insight, we propose two Specialist-MLLM collaboration strategies to enhance REC performance. The first strategy, Slow-Fast Adaptation (SFA), introduces an adaptive task routing mechanism that dynamically assigns simpler, detection-oriented tasks to lightweight Specialist Models, while delegating complex tasks requiring compositional reasoning to MLLMs. This adaptive and flexible collaboration not only improves efficiency but, more importantly, enhances overall REC performance. Additionally, we identify common error patterns in both types of models, particularly their tendency to mistakenly identify objects not belonging to the described category. To mitigate this issue, we propose a simple yet effective target-refocus strategy, which obviously reduces localization confusion for both Specialist Models and MLLMs. Notably, the SFA strategy is entirely training-free, making it highly practical for real-world applications. The second strategy, Candidate Region Selection (CRS), aims to improve recall and precision by generating multiple object candidates using Specialist Models, followed by leveraging MLLMs’ reasoning capabilities to select the correct target. While MLLMs inherently possess reasoning capacity, they are not explicitly trained to approach selection tasks as generative problems, limiting their effectiveness in such scenarios. To address this, we conduct instruction tuning on external datasets, ensuring that MLLMs can effectively address selection tasks while avoiding knowledge leakage. This instruction tuning process generalizes well across various REC datasets.
@@
The preliminary version of this work was published in [7]. In this paper, we made significant extensions. While the FineCops-Ref benchmark was introduced in [7], this work goes further by proposing two Specialist-MLLM collaboration strategies to effectively tackle the fine-grained compositional REC task, both within FineCops-Ref and across other external benchmarks. Additionally, we present more comprehensive experiments and analyses, providing a more self-contained and in-depth exploration of the fine-grained compositional REC task.
@@ II-A Benchmarking Referring Expression Comprehension
Referring Expression Comprehension (REC) is a vision-language task where models must localize a target object in an image based on a natural language referring expression, requiring compositional understanding of attributes (e.g., color, size), spatial relationships, and object categories to disambiguate the referent from other entities. While early benchmarks like RefCOCO/+/g [2, 3, 4] established foundational evaluation protocols, subsequent analyses revealed critical limitations. For instance, [5] and [6] demonstrated that models could exploit dataset biases rather than genuinely parse linguistic structure: [6] found that up to 83.7% of RefCOCOg test instances could be solved via keyword matching or visual feature dominance, bypassing compositional reasoning. To address this, [6] introduced Ref-Adv, a dataset where perturbed expressions refer to alternate objects, forcing models to engage with linguistic structure. Subsequent efforts have prioritized compositional reasoning. CLEVR-Ref+ [8] is a synthetic dataset emphasizing relationships, attributes, and linguistic logic. Cops-Ref [9] and Ref-Reasoning [10] use GQA scene graphs [11] and rule-based methods to create large-scale compositional referring expression comprehension datasets in real-world scenarios. Cops-Ref additionally introduces distracting images based on attributes, relationships, and target names. Recent benchmarks like GITM-MR [12] explores mismatched relationship in the REC task. ARPGrounding [13] evaluates the model’s compositional reasoning ability by constructing referring expression pairs that target distinct objects of the same category within a single image, differentiated solely through their attributes or relational contexts. RefEgo [14] and OmniLabel [15] consider out-of-distribution scenarios where referred targets do not exist in the image.
Building on these advances, we propose a new REC benchmark that more comprehensively assesses the compositional reasoning abilities of multimodal models. Our dataset incorporates controlled difficulty levels, compelling MLLMs to reason across object categories, attributes, and multi-hop relationships. Crucially, we also introduce negative text and image samples to evaluate model resilience against misalignments and hallucinations, providing a more thorough assessment of their true visual grounding capabilities. By simultaneously addressing compositional reasoning and incorporating negative samples, FineCops-Ref establishes a more realistic and challenging benchmark for advancing REC research in the MLLM era.
@@ II-B Specialist REC Models
Figure 1: The data construction pipeline of FineCops-Ref. Given an image, we first generate paths based on its scene graph. Then, we fill paths into templates and obtain the positive referring expression through LLM rewriting. Meanwhile, we utilize LLM to generate negative expressions, and based on this, we employ diffusion model to create fine-grained editing negative images.
@@ III FineCops-Ref Benchmark
In this section, we provide a detailed description of the FineCops-Ref dataset construction. Our dataset comprises both positive and negative samples. Positive samples consist of paired images and textual expressions, while negative samples contain unpaired images and textual expressions. The negative samples are specifically designed to evaluate the model’s ability to reject instances where the target object is absent from the image. The dataset construction pipeline is illustrated in Fig. 1.
III-A Creating Positive Data
Path Generation. We employ image scene graphs from GQA [11] for path generation. The scene graphs contain detailed information about objects, attributes, and relations. To ensure accuracy, we first filter the objects based on their suitability as target or related objects. We leverage annotations from InstInpaint [54] and apply additional filters such as keywords and object size. Next, we generate several paths for each of the filtered objects, as show in Fig. 1(a). To eliminate ambiguity, we utilize unique attributes or relations to identify the target object that shares the same category as other objects in the image to ensure that every generated path is unique.
Data Categorization. We categorize the paths into three difficulty levels based on the complexity of fine-grained reasoning. Level 1 indicates that there are no other objects in the image belonging to the same category as the target object, such as the TV in Fig. 1. In this case, the model can locate the target without requiring contextual understanding. Level 2 signifies the presence of another object with the same category as the target in the image, where the target can be distinguished through one unique attribute or relation. Level 3 requires at least two or more relationships and attribute information. The difficulty levels are determined by the intricacy of fine-grained reasoning, rather than the complexity of the textual description.
@@ III-A Creating Positive Data
Expression Generation. To mitigate potential biases in the scene graphs, we first apply frequency-based sampling of relationships, attributes, and object categories along the generated paths. Subsequently, we use predefined templates to generate referring expressions. The details of predefined templates are provided in Appendix.
To further augment the naturalness and diversity of these expressions, we leverage GPT-3.5-turbo to rewrite the referring expressions. By incorporating well-designed instructions and examples, we achieve a wider range of linguistically diverse and natural expressions. The prompts used for rewriting are listed in Appendix.
Human Filter. Due to inherent limitations in the scene graph annotations, the generated paths, particularly for Level 2 and Level 3, may contain inaccuracies, leading to non-unique target references. To address this, human annotators manually filtered the test set. Further details can be found in Appendix.
III-B Generating Negative Data
To conduct a thorough and systematic assessment of REC of existing VLMs, we generate hard negatives from both textual and visual sources. Similar to positive data, negative data are categorized into different levels based on difficulty. Negative Level 1 involves alterations to the target object in the expressions, which are relatively straightforward for the model to identify. Level 2 involves modifications to the related objects, disrupting the contextual information and posing a greater challenge for existing models to recognize.
Generating Negative Expressions. Our set of negative expressions encompasses a wide range of challenging types. Inspired by CREPE [55] and SUGARCREPE [56], we consider various forms of hard negatives. In total, FineCops-Ref covers 5 fine-grained types of hard negative expressions. These types can be broadly classified into two categories: Replace and Swap. The Replace category involves generating negative expressions by substituting a portion of the original expression, whether it is an object, attribute, or relation. We utilize LLM to determine the most appropriate negative word, ensuring that the negative expression is genuinely negative while only slightly deviating from the original expression. We experimented with various replacement methods and found that LLM-based replacements performed the best, as discussed in Appendix. The Swap category entails generating negative expressions by interchanging two attributes or objects within the same category. We further employ LLM to rewrite these new expressions.
Generating Negative Images. We consider the necessity of negative images from the following aspects. First, negative images enables a more thorough assessment of models’ visual parsing capabilities. Additionally, evaluations conducted by Visualgptscore [57] suggest that negative expressions may lack plausibility and fluency and can be detected by language prior.
We generate hard negative images with subtle differences from the original, such as modifications to objects, attributes, or relations. For simple positional relationships, we employ horizontal flips. For more intricate modifications involving objects and attributes, we utilize PowerPaint [58], an exceptional image inpainting model offering versatility, to perform precise edits on the images. To guide PowerPaint in editing the image, we utilize LLM-generated replacements as textual guides and the bounding boxes as masks. Overall, FineCops-Ref encompasses 5 distinct types of challenging negative images. Further details on the data types, along with example expressions and images, can be found in Appendix.
Negative Data Debiasing. During the generation of negative samples, some implausible or incoherent expressions, as well as unreasonable and easily distinguishable negative images, are inevitable. We employed several techniques to filter out unsuitable samples and improve the quality of the benchmark. For negative expressions, we employ the Adversarial Refinement technique proposed by SUGARCREPE [56], which helps mitigate biases and unintended artifacts in the dataset.
To exclude inappropriate and excessively unreasonable negative images, we employ a multi-step filtering process. First, we use CLIP [59] to ensure that the similarity between the negative text and the positive image is lower than the similarity between the positive text and the positive image. Next, we apply the diffusion-generated inspection model DIRE [60] to filter out excessively unnatural images, excluding those with scores above 0.2. Finally, we use DINOv2 [61] to compute the image-image similarity between the positive image and the generated negative images, retaining the candidate negative image with the highest DINOv2 score from a set of 10 candidates, thereby minimizing noise.
@@ VI Conclusion
In this work, we introduced a high-quality dataset, FineCops-Ref, specifically designed for fine-grained compositional Referring Expression Comprehension (REC). This dataset incorporates varying difficulty levels and includes negative samples, providing a comprehensive evaluation of REC models’ capabilities in object detection, cross-modal reasoning, and rejection of non-existent objects—key aspects for real-world REC applications. To address these challenges, we proposed a Specialist-MLLM collaboration strategy that leverages the strengths of both object detection and cross-modal reasoning. This strategy is implemented through two approaches: Slow-Fast Adaptation (SFA), which dynamically assigns tasks to the appropriate specialist or MLLM model based on task difficulty, and Candidate Region Selection (CRS), which generates multiple region candidates using specialist models and selects the target through a multi-choice strategy. Extensive experiments on the FineCops-Ref dataset, along with two other challenging REC datasets, demonstrate the effectiveness of our proposed methods.

### v3

@@ Abstract
Referring Expression Comprehension (REC) is a foundational cross-modal task that evaluates the interplay of language understanding, image comprehension, and language-to-image grounding. It serves as an essential testing ground for Multimodal Large Language Models (MLLMs). To advance this field, we introduced a new REC dataset in our previous conference paper, characterized by two key features. First, it is designed with controllable difficulty levels, requiring multi-level fine-grained reasoning across object categories, attributes, and multi-hop relationships. Second, it incorporates negative text and images generated through fine-grained editing and augmentation, explicitly testing a model’s ability to reject scenarios where the target object is absent—an often-overlooked yet critical challenge in existing datasets. In this extended work, we propose two new methods to tackle the challenges of fine-grained REC by combining the strengths of Specialist Models and MLLMs. The first method adaptively assigns simple cases to faster, lightweight models and reserves complex ones for powerful MLLMs, balancing accuracy and efficiency. The second method lets a specialist generate a set of possible object regions, and the MLLM selects the most plausible one using its reasoning ability. These collaborative strategies lead to significant improvements on our dataset and other challenging benchmarks. Our results show that combining specialized and general-purpose models offers a practical path toward solving complex real-world vision-language tasks. Our dataset and code are available at https://github.com/sleepyshep/FineCops-Ref.
@@ I Introduction
To address these limitations in existing datasets, we introduced the FineCops-Ref benchmark in our previous work [7]. FineCops-Ref is specifically designed to enhance fine-grained reasoning in REC by incorporating controlled difficulty levels and meticulously constructed negative samples. The difficulty levels compel MLLMs to reason across object categories, attributes, and multi-hop relationships, categorized based on the number of attributes and relationships necessary for accurate localization. The inclusion of negative samples evaluates model resilience against misalignments and hallucinations, providing a more rigorous assessment of their true visual grounding capabilities. By addressing both compositional reasoning and negative sample inclusion, FineCops-Ref establishes a more realistic and challenging benchmark for advancing REC research in the MLLM era.
@@
Extensive evaluations on the FineCops-Ref benchmark also reveal a critical gap in current vision-language models [7]: while MLLMs demonstrate strong generalization and reasoning abilities, they often struggle with fine-grained localization and compositional grounding, particularly when precise object differentiation is required. In contrast, Specialist Models, which are typically lightweight and trained for specific vision tasks, excel in efficient and accurate low-level perception but lack the broad contextual reasoning capabilities of MLLMs. Given their complementary strengths, an effective solution is to leverage both models strategically—assigning simpler, perception-driven tasks to Specialist Models while reserving MLLMs for complex reasoning. This collaborative approach not only enhances accuracy and robustness but may also improve computational efficiency, as MLLMs are invoked only when necessary.
@@
Building on this insight, this extended work introduces two collaboration strategies between Specialist Models and MLLMs to improve REC performance. The first strategy, Slow-Fast Adaptation (SFA), introduces an adaptive task routing mechanism that dynamically assigns simpler, detection-oriented tasks to lightweight Specialist Models, while delegating complex tasks requiring compositional reasoning to MLLMs. This adaptive and flexible collaboration not only improves efficiency but, more importantly, enhances overall REC performance. Additionally, we identify common error patterns in both types of models, particularly their tendency to mistakenly identify objects not belonging to the described category. To mitigate this issue, we propose a simple yet effective target-refocus strategy, which obviously reduces localization confusion for both Specialist Models and MLLMs. Notably, the SFA strategy is entirely training-free, making it highly practical for real-world applications. The second strategy, Candidate Region Selection (CRS), aims to improve recall and precision by generating multiple object candidates using Specialist Models, followed by leveraging MLLMs’ reasoning capabilities to select the correct target. While MLLMs inherently possess reasoning capacity, they are not explicitly trained to approach selection tasks as generative problems, limiting their effectiveness in such scenarios. To address this, we conduct instruction tuning on external datasets, ensuring that MLLMs can effectively address selection tasks while avoiding knowledge leakage. This instruction tuning process generalizes well across various REC datasets.
@@
This work extends our previous conference paper [7], which introduced the FineCops-Ref benchmark. In this journal version, we make three major contributions: (1) we propose two novel Specialist–MLLM collaboration strategies tailored for fine-grained compositional REC; (2) we conduct more extensive experiments across FineCops-Ref and multiple external benchmarks; and (3) we provide deeper analyses and ablations, offering a more self-contained and comprehensive exploration of the task.
@@ II-A Benchmarking Referring Expression Comprehension
Referring Expression Comprehension (REC) is a vision-language task where models must localize a target object in an image based on a natural language referring expression, requiring compositional understanding of attributes (e.g., color, size), spatial relationships, and object categories to disambiguate the referent from other entities. While early benchmarks like RefCOCO/+/g [2, 3, 4] established foundational evaluation protocols, subsequent analyses revealed critical limitations. For instance, [5] and [6] demonstrated that models could exploit dataset biases rather than genuinely parse linguistic structure: [6] found that up to 83.7% of RefCOCOg test instances could be solved via keyword matching or visual feature dominance, bypassing compositional reasoning. To address this, [6] introduced Ref-Adv, a dataset where perturbed expressions refer to alternate objects, forcing models to engage with linguistic structure. Subsequent efforts have prioritized compositional reasoning. CLEVR-Ref+ [8] is a synthetic dataset emphasizing relationships, attributes, and linguistic logic. Cops-Ref [9] and Ref-Reasoning [10] use GQA scene graphs [11] and rule-based methods to create large-scale compositional referring expression comprehension datasets in real-world scenarios. Cops-Ref additionally introduces distracting images based on attributes, relationships, and target names. Recent benchmarks like GITM-MR [12] explores mismatched relationship in the REC task, while ARPGrounding [13] evaluates the model’s compositional reasoning ability by constructing referring expression pairs that target distinct objects of the same category within a single image, differentiated solely through their attributes or relational contexts. RefEgo [14] and OmniLabel [15] consider out-of-distribution scenarios where the referred targets do not exist in the image.
The challenges in rigorously evaluating compositional reasoning are not unique to REC. Across various related vision-language tasks, such as image captioning, image-to-text retrieval, and visual question answering, recent evaluations similarly reveal that current vision-Language Models struggle with compositional understanding [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. For instance, prominent benchmarks like FOIL-it [16], VALSE [17], and COCO-Counterfactuals [20] are specifically designed to test compositional understanding by creating “hard negative” image-caption pairs that highlight subtle distinctions. These broader evaluation efforts underscore the critical need for benchmarks that can truly test deep linguistic-visual understanding and resilience to subtle distractors in complex, compositional scenarios – a requirement particularly pronounced in fine-grained REC where minor linguistic cues can drastically change the target.
Building on these advances, we introduced FineCops-Ref in our previous work [7], a new REC benchmark specifically designed to more comprehensively assess the fine-grained compositional reasoning abilities of multimodal models. Based on this challenging benchmark, we proposes novel Specialist-MLLM collaboration strategies and provides more extensive experiments to further advance fine-grained compositional REC.
@@ II-B Specialist REC Models
Figure 1: The data construction pipeline of FineCops-Ref [7]. Given an image, we first generate paths based on its scene graph. Then, we fill paths into templates and obtain the positive referring expression through LLM rewriting. Meanwhile, we utilize LLM to generate negative expressions, and based on this, we employ diffusion model to create fine-grained editing negative images.
@@ III FineCops-Ref Benchmark
In this section, we provide a summary of the FineCops-Ref dataset introduced in our previous work [7]. FineCops-Ref is a benchmark specifically designed to address the limitations of existing datasets in evaluating fine-grained compositional reasoning and robustness to negative samples in Referring Expression Comprehension. It comprises both positive and negative samples, carefully constructed to provide a more challenging and realistic evaluation setting for Multimodal models. The dataset construction pipeline is illustrated in Fig. 1. The detailed methodology for constructing FineCops-Ref, including path generation, data categorization, expression generation, human filtering, and the generation of negative data (both expressions and images), was introduced in our previous conference work [7].
III-A Key Characteristics
FineCops-Ref is distinguished by the following key characteristics, which make it a suitable and challenging benchmark for evaluating advanced REC models and our proposed collaboration strategies:
Controlled Difficulty Levels: The dataset incorporates three controlled difficulty levels (Level 1, 2, and 3) based on the complexity of fine-grained reasoning required to uniquely identify the target object. This necessitates models to reason across object categories, attributes, and multi-hop relationships. Level 1 involves simple localization without distractors of the same category, Level 2 requires distinguishing the target using a single attribute or relation among same-category distractors, and Level 3 demands the interpretation of two or more relationships and attribute dependencies.
Comprehensive Negative Samples: FineCops-Ref includes both negative expressions and negative images. These negative samples are specifically designed as “hard negatives” with subtle differences from positive instances, evaluating the model’s ability to correctly reject instances where the described object is absent or the expression/image is subtly mismatched. The construction involves various types of challenging negative samples generated through careful editing and perturbation.
@@ VI Conclusion
In this work, we build upon our previously introduced FineCops-Ref dataset, which was designed for fine-grained compositional Referring Expression Comprehension (REC). FineCops-Ref features controllable difficulty levels and carefully constructed negative samples, enabling comprehensive evaluation of REC models in object detection, cross-modal reasoning, and rejection of non-existent targets—capabilities critical for real-world applications. To further advance this task, we proposed a Specialist-MLLM collaboration strategy that leverages the strengths of both object detection and cross-modal reasoning. This strategy is implemented through two approaches: Slow-Fast Adaptation (SFA), which dynamically assigns tasks to the appropriate specialist or MLLM model based on task difficulty, and Candidate Region Selection (CRS), which generates multiple region candidates using specialist models and selects the target through a multi-choice reasoning strategy. Extensive experiments on the FineCops-Ref dataset, along with two other challenging REC datasets, demonstrate the effectiveness of our proposed methods.